-4x+y=10 3x+19y=32 solve using substitition method
Found 2 solutions by Cromlix, MathTherapy:
Answer by Cromlix(4381)

-4x + y = 10.........1
3x + 19y = 32.......2
Multiply (1) by 3 and (2)by 4
-12x + 3y = 30.......3
12x + 76y = 128.....4
Add (3) + (4)
79y = 158
y = 2
Substitute y = 2 into (2)
3x + 19y = 32
3x + 19(2) = 32
3x + 38 = 32
3x = 32 - 38
3x = -6
x = -2
x = -2 y = 2.
Hope this helps.
:-)

Answer by MathTherapy(10555)


-4x+y=10 3x+19y=32 solve using substitition method

- 4x + y = 10 ----- eq (i)

3x + 19y = 32 ------ eq (ii)

Solve eq (i) for y and substitute that value for y in eq (ii). This will give you the value of x. Substitute the x-value into any of the 2 original equations [eq (i) or eq (ii)] to determine the y-value.